OpenStudy (anonymous):

*I WILL GIVE MEDAL* When the sum of 1.9 and 2.2 is subtracted from the product of 1.9 and 2.2, what is the difference?? @TheSmartOne @amorfide @kittenlover731

OpenStudy (anonymous):

ok so the sum of 1.9 and 2.2 means you need to add the them together so do 1.9+2.2= ?

OpenStudy (anonymous):

yep, now subtract the product (4.1) from the product of 1.9 and 2.2. Product means to multiply so do that 1.9*2.2= ?

OpenStudy (anonymous):

4.18 @kittenlover731

OpenStudy (anonymous):

yep! now the last part says "what is the difference?" difference means subtraction so subtract your two products 4.18 - 4.1= ?

OpenStudy (anonymous):

.08 @ki

OpenStudy (anonymous):

@kittenlover731